In a G.P., the product of first four terms is 4 and the second term is the reciprocal of the fourth term. Show that the sum of infinite terms of the G.P. is -8, -8/3, 8/3, 8 

Grade:11

2 Answers

iit jee
14 years ago

let 4 nos be a/r2,a/r,ar,ar2.

a= 4

a = 21/2

ar2 = r/a

r = 1/2

sum of infinite terms of GP = a/1-r = 8 

Vaishnav
5 years ago
Selection the terms of gp=a/r^2,a/s,as,ar^2
Product is your numbers will be a^4=4
a=2^1/2
ar^2=r/a
r=1/2
Sum of infinite gp will be =a/1-r=8